A converging lens of focal length 0.40cm is combined with a diverging lens of focal length 0.20cm the power of combination of the lens is


1
Expert's answer
2021-04-19T17:11:09-0400

Focal length of the converging lens is 0.40 cm i.e. f1 ​= 0.004 m

Focal length of the diverging lens is 0.20 cm i.e. f2​ = −0.002 m

"P=\frac{1}{f}=\frac{1}{f_1}+\frac{1}{f_2}=\frac{1}{0.004}+\frac{1}{-0.002}=-250" D